OpenStudy (anonymous):

what is the general term of this sequence: -1, 2, -3, 4, -5

OpenStudy (campbell_st):

is something like \[-1 \times ( - \frac{n +1}{n})\] where n = 1, 2, 3, ...... given a 1st term of -1

OpenStudy (campbell_st):

that generates the terms -2, 3, -4, 5 etc.

OpenStudy (anonymous):

isn't that wrong? for n=2 it is \[-1\times \left( \frac{ 2+1 }{ 2 } \right)\] which is 3/2

OpenStudy (campbell_st):

got it \[T_{n} = (-1)^n \times n\]

OpenStudy (campbell_st):

lol... I looked for to much in the question... sorry

OpenStudy (anonymous):

ah i found it out. it should be \[(-1^n)*-((n+1)/1^n)\]

OpenStudy (campbell_st):

that doesn;t work for n = 1 T1 = (-1^1)* -1(1 +1)/1^1 T1 = 2 you need T1 = -1

OpenStudy (anonymous):

yeah i see that now. but the last one you gave me works. thanks :)

OpenStudy (campbell_st):

lol.. is was simpler than it looked..

OpenStudy (anonymous):

yeah it was